OpenStudy (anonymous):

The monthly payment for a car lease is $530 for 18 months. The down payment is $2,000. The residual value is $7,000. What is the total cost of the lease?

OpenStudy (anonymous):

Total cost = down payment + 18 * 530 = $11,540

OpenStudy (anonymous):

Total lease = 2,000 + 18 * 530 = $11,540